# Trials Factor for Semi-Supervised NN Classifiers in Searches for Narrow Resonances at the LHC

### v1

@@ 1 Introduction
The discovery of the Brout-Englert-Higgs boson [1, 2, 3, 4], a scalar resonance, by the ATLAS [5] and CMS [6] collaborations in 2012, marked a significant milestone, completing the particle spectrum of the Standard Model (SM). The endeavour to unearth beyond the SM (BSM) physics hinges significantly on the searches for new particles in the form of narrow resonances. However, the lack of confirmed signals in inclusive or benchmark analyses implies that potential new resonances concealed within the data are likely to be subtle. Such subtlety could be the result of complex production and decay topologies that evade coverage by standard analyses (for a review of the subject see Ref. [7]). Consequently, BSM resonances might have evaded detection, lurking unnoticed within uncharted or inadequately explored regions of the phase space.
@@
Machine Learning has the substantial promise of addressing this challenge. It has already demonstrated considerable effectiveness in probing areas of special interest, exemplified by the detection of the SM Higgs boson. These “standard” fully supervised techniques, which rely on the knowledge of event-by-event truth-level information, called “labels", can falter if the manifestation of BSM physics diverges from the model used as a reference. This means that if the simulated data on which these models are trained is too specific or flawed, due to bias or insufficient truth-level information, the model will learn these artifacts of the simulation and might not recognise inherent signals. To reduce this model dependency, weakly supervised methods can be used [8, 9, 10] which do not enforce labels for the signal events.
@@
The classification without labels (CWoLa) study, presented by the ATLAS collaboration, introduces weak learning techniques to reduce the dependency of the recognition mechanism on simulated data [11]. The study uses mixed samples (combining signal and background events in each sample) to train binary classification models. It reveals that weakly supervised models can classify signal events with success comparable to the fully supervised method [12, 13, 14, 15]. Variations of the weakly supervised method have been successfully implemented for resonance detection in Refs. [16, 17, 18, 19]. A further advantage of the weakly supervised method is shown to be its ability to be trained on impure  [^1] An impure sample refers to a dataset containing not only the signal and background of interest, but also additional signals. mixed samples [20], or directly on data, reducing the need for simulation. Research on the use of weak supervision in particle physics has been extended to include multi-model classifiers [21], graph neural networks [22], with the potential inclusion of transfer learning [23]. Finally, the viability of the methodologies for narrow resonance searches was studied in Refs. [17, 16] and a review of weak supervision was presented in Ref. [24].
@@
This weakly or semi-supervised method can be adapted to use partial labelling, with a labelled background sample and an unlabelled mixed (signal and background) sample. This methodology allows the classifier to learn to distinguish background events from signal events by being trained on a background-only sample of simulated events. By subtracting the learned background events from the mixed sample, one is left only with the signal events of interest. The mixed sample can therefore consist of actual LHC data, allowing for a model-independent extraction of signal events. This semi-supervised technique, implemented via a side-band analysis, is therefore able to reduce model dependency and detect BSM physics with topological variations between the possible signal and side-band regions [25].
@@
In general, model-independent searches for narrow resonances are of great interest. In recent years the field of particle physics has experienced a growing plethora of anomalous experimental results (for recent reviews see Refs. [27, 28]). Many of them are statistically significant, continue to grow, and remain unexplained by state-of-the-art SM calculations. While some anomalies might eventually find resolution within the SM, persisting ones could contain signs of new physics and may serve as a guide to model building and experimental searches. Of particular relevance to this paper are the multi-lepton anomalies at the LHC [29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. These comprise excesses in final states with charged leptons (electrons and muons), missing energy (with and without hadronic and $b$-jets), in processes that have SM Higgs-like topologies. They can be explained by introducing (at least) three additional scalar fields, $H$ and $S$, and $S^{\prime}$ [39, 40]. The scalar $H$ has a mass in the ballpark of 270 GeV, while $S(S^{\prime})$ have masses close to that of the SM Higgs. From the di-lepton component of the multi-lepton anomalies, and assuming $S\rightarrow W^{+}W^{-}\rightarrow\ell\ell\nu\nu,\ell=e,\mu$, the mass of $S$ was determined to be $m_{S}=150\pm 5$ GeV [31], which was later confirmed in Refs. [33, 38]. The first indication of a narrow resonance within this mass range of the di-photon spectrum was reported at $m_{S}=151.5$ GeV in Ref. [41] and further corroborated with more data in Ref. [42]. [^2] At the time of the manuscript submission new excesses at the same mass have been reported in Ref. [43]. Furthermore, an excess around $95$ GeV, denoted $S^{\prime}$ (which acts as the main source of $b$ quarks), has been observed [44, 45, 46, 47, 37, 42]. This provides a compelling reason to seek Higgs-like resonances near the electroweak (EW) scale, an area where SM backgrounds are quite significant. The use of topological constraints through semi-supervision could be crucial in this context, as they help to dampen the impact of these backgrounds, while reducing model dependencies in the search.
@@ 2.1 Monte Carlo Simulation
The dominant background, for our example search for the $S(150\,{\rm GeV})\rightarrow Z\gamma$ decay, is the production of a prompt photon and a $Z$ boson [48], accounting for more than $92\%$ of the total background [49]. [^3] A secondary contribution comes from the production of $Z$ bosons in association with jets ($Z+$jets), with one jet misidentified as a photon. The $Z+$jets background is not considered in this study as its contribution is derived from real data at the LHC. A sample of $000110^{\prime}$ 000 SM $Z\gamma$ MC events are simulated using MadGraph5 aMC@NLO 2.6.7 with Next-to-Leading Order accuracy in QCD [50]. The parton showering and hadronisation were simulated with PYTHIA 8.2 [51] using ATLAS A14 event tuning and NNPDF2.3 next-to-leading order simulation with a leading-order parton distribution function set[52]. Events were processed with Delphes 3 [53], which provides an approximate fast simulation of the ATLAS detector. A set of generator-level cuts were applied, where the transverse momentum of the photons was required to be greater than $25$ GeV. Hadronic jets were reconstructed using the anti-$k_{t}$ algorithm [54] with the radius parameter, $R=0.4$, as implemented in the FastJet 3.2.2 [55] package, and jets with $p_{\textrm{T}}>30$ GeV and $|\eta|<4.7$ are considered. In addition, jets originating from bottom quarks are identified as $b$-jets with $b$-tagging algorithm [56]. Reconstructed jets overlapping with photons, electrons or muons in a cone of size $R=0.4$ are removed. Electrons and muons are required to have $p_{\textrm{T}}>25$ GeV and $|\eta|<2.5$.
The observables to be used as kinematic features in our study are:
@@
  • The number of jets ($N_{j}$), and the number of central jets ($N_{cj}$).
The feature distributions of the MC dataset are shown in Figure 1. All observables, excluding $m_{\ell\ell\gamma}$, are used as inputs for the NN classifier. The $m_{\ell\ell\gamma}$ distribution, and any features with high correlation with it, cannot be used to train the NN due to the fact that the training samples, side-band and mass window, are defined on this mass.
@@ 3 Methodology
Figure 4: Diagram illustrating the methodology for a pseudo-experiment to assess possible trails factors when using a semi-supervised NN in the search for a narrow resonance. The process starts with the generation of a sample dataset, followed by the training of the semi-supervised NN. Subsequent stages include a background rejection scan to extract local samples from the NN response, fitting signal and background models to each local sample, and finally the computation of the signal significance.
@@ 3.3 Significance calculation for a resonance in the invariant mass distributions
In this analysis, the resonant anomaly detection search strategy is used to discover evidence for BSM physics. The new physics signal is expected to originate from a decay of a new particle, $S$, in the form of a resonance in the corresponding invariant mass. The investigation therefore involves searching for an excess with respect to the continuous SM background. Since we want to examine the probability of a fake signal, we define a signal region (or mass window) in the invariant mass, $m_{\ell\ell\gamma}$, where an excess of signal events can be observed. We select a fixed center of mass of $m_{\ell\ell\gamma}\approx 150$ GeV, motivated by the corresponding excess [31, 38, 41, 42]. The background (or side-band) region is defined as the mass range surrounding the signal region where only SM background events are expected. The signal (mass-window) range is defined as $144\leq m_{\ell\ell\gamma}\leq 156$ GeV and the background (side-band) range as $168132<m_{\ell\ell\gamma}<$ 168 GeV excluding the mass-window.
@@
where $c_{0}$, $c_{1}$ and $c_{2}$ are fit parameters and $\epsilon_{m}$ is the invariant mass, $m_{\ell\ell\gamma}$, divided by the center of mass energy, i.e. $13$ TeV. Excesses of signal events within the mass window are captured using a Gaussian function, Equation 2:
$g(m)=c_{s}\cdot e^{-\frac{(m-\mu_{s})^{2}}{2\cdot\sigma_{s}^{2}}}$ (2)
where $m$ is the invariant mass, $c_{s}$ is the parameter measuring the height of the peak, $\mu_{s}$ is the position of the center of the peak, and $\sigma_{s}$ is the search resolution. In this analysis, the $\mu_{s}$ is fixed at $150$ GeV and $\sigma_{s}$ is set to the search resolution of $2.4$ GeV [49].
@@
where $N_{\rm s}$ and $N_{\rm b}$ are the number of signal and background events respectively. Each pseudo-experiment will therefore produce six local significance values, reflecting each background rejection sample. An example of the signal and background fits for a single iteration of the pseudo-experiment are shown in Appendix A, Figure 8.
@@ 4 Results and Discussion
This methodology, similar to mass scans in high-energy physics, segments our analysis through the use of background rejection (BR) samples, where each threshold acts as a local category. Each category, however, is not statistically independent as the events in each sample overlap. This approach facilitates an examination of the impact of background rejection levels on our results, enabling a nuanced understanding of their influence on our findings. The probability density functions (PDF) of the significances for the six different background rejection cuts ($f(\sigma)_{\rm BR}$) are shown in Figure 4(a). Note that these distributions have approximately a Gaussian shape, however with a standard deviation between $0.65$ and $0.71$. This decrease in variance w.r.t. the standard Gaussian is due to the constraints within the resonance fit imposed a priori via the shape of the resonance and the side-bands definitions. Note that the peak of the PDFs shifts to the right with an increasing background rejection cut. Such a shift emerges from the interplay between signal and background modelling, amplified by the diminishing event counts in samples subjected to higher background rejection cuts.
@@
where BR is a label indicating the local background rejection cut. We then compute the probability that, for each background reject cut, a significance of up-to a value $\sigma^{T}=1\sigma$, $2\sigma$ and $3\sigma$ is obtained via the cumulative distribution function (CDF):
@@
The “global” PDF
$f(\sigma)_{\rm global}=\sum_{i=1}^{6}f(\sigma)_{i},$ (6)
where $i$ represents the different background rejection cuts, is used to the average of all background rejection categories. It is shown in Figure 6 from where one can see that it approximates a Gaussian distribution with a standard deviation of $0.69$ centred around a mean of $0.15$. The corresponding global $p$-value is calculated as
@@
Figure 6: Probability density distribution of the significance from the global sample, i.e. averaged over the different background rejection cuts.
@@
Figure 7: The ratio $p_{\rm X}/p_{0\%\,{\rm BR}}$ for $X=10\%,20\%,30\%,40\%,50\%$ BR cuts and $p_{\rm global}/p_{0\%\,{\rm BR}}$, corresponding to the trails factor, as a function of the significance threshold, $\sigma^{T}$.

### v2

@@ 1 Introduction
The discovery of the Brout-Englert-Higgs boson [1, 2, 3], a scalar resonance, by the ATLAS [4] and CMS [5] collaborations in 2012, marked a significant milestone, completing the particle spectrum of the Standard Model (SM). The endeavour to unearth beyond the SM (BSM) physics hinges significantly on the searches for new particles in the form of narrow resonances. However, the lack of confirmed signals in inclusive or benchmark analyses implies that potential new resonances concealed within the data are likely to be subtle. Such subtlety could be the result of complex production and decay topologies that evade coverage by standard analyses (for a review of the subject see Ref. [6]). Consequently, BSM resonances might have evaded detection, lurking unnoticed within uncharted or inadequately explored regions of the phase space.
@@
Machine Learning has the substantial promise of addressing this challenge. It has already demonstrated considerable effectiveness in probing areas of special interest, exemplified by the detection of the SM Higgs boson. These “standard” fully supervised techniques, which rely on the knowledge of event-by-event truth-level information, called “labels", can falter if the manifestation of BSM physics diverges from the model used as a reference. This means that if the simulated data on which these models are trained is too specific or flawed, due to bias or insufficient truth-level information, the model will learn these artifacts of the simulation and might not recognise inherent signals. To address this challenge, model-independent searches utilise unsupervised [7, 8, 9], semi-supervised [10, 11] or weakly supervised [12, 13] techniques. Unsupervised techniques do not rely on any labels for events, thus avoiding any dependency on a model. In contrast, semi-supervised techniques apply labels only to background events, enabling a model-independent extraction of signal events. Weakly supervised techniques, on the other hand, utilize the proportions of signal to background events in the training samples, limiting the model dependence by relying on partial labeling. These techniques are therefore most often used to look for deviations from the background-only hypothesis.
@@
The classification without labels (CWoLa) study, presented by the ATLAS collaboration, introduces weak learning techniques to reduce the dependency of the recognition mechanism on simulated data [14]. The study uses mixed samples (combining signal and background events in each sample) to train binary classification models. It reveals that weakly supervised models can classify signal events with success comparable to the fully supervised method [15, 16, 17, 18]. Variations of the weakly supervised method have been successfully implemented for resonance detection in Refs. [19, 20, 21]. An advantage of the model agnostic, weakly and semi- supervised, methods is the ability to be trained on impure  [^1] An impure sample refers to a dataset containing not only the signal and background of interest, but also additional signals. mixed samples [22], or directly on data, reducing the need for simulation. Research on the use in particle physics has been extended to include multi-model classifiers [23], graph neural networks [24], with the potential inclusion of transfer learning [25]. Finally, the viability of the model-independent methodologies for narrow resonance searches was studied in Refs. [20, 19] and a review of weak supervision was presented in Ref. [26].
@@
This weakly-supervised method can be adapted to use a labeled background sample containing no signal events, alongside an unlabeled mixed sample containing both signal and background events. This methodology allows the classifier to learn to distinguish background events from signal events by being trained on a background-only sample of simulated events. By subtracting the learned background events from the mixed sample, one is left only with the signal events of interest. The mixed sample can therefore consist of actual LHC data, allowing for a model-independent extraction of signal events. This semi-supervised technique, implemented via a side-band analysis, is therefore an extension of the weak supervision method ideal for reducing model dependency and detect BSM physics with topological variations between the possible signal and side-band regions [27].
@@
In general, model-independent searches for narrow resonances are of great interest. In recent years the field of particle physics has experienced a growing plethora of anomalous experimental results (for recent reviews see Refs. [29, 30]). Many of them are statistically significant, continue to grow, and remain unexplained by state-of-the-art SM calculations. While some anomalies might eventually find resolution within the SM, persisting ones could contain signs of new physics and may serve as a guide to model building and experimental searches. Of particular relevance to this paper are the multi-lepton anomalies at the LHC [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. These comprise excesses in final states with charged leptons (electrons and muons), missing energy (with and without hadronic and $b$-jets), in processes that have SM Higgs-like topologies. They can be explained by introducing (at least) three additional scalar fields, $H$ and $S$, and $S^{\prime}$ [37, 46]. The scalar $H$ has a mass in the ballpark of 270 GeV, while $S(S^{\prime})$ have masses close to that of the SM Higgs. From the di-lepton component of the multi-lepton anomalies, and assuming $S\rightarrow W^{+}W^{-}\rightarrow\ell\ell\nu\nu,\ell=e,\mu$, the mass of $S$ was determined to be $m_{S}=150\pm 5$ GeV [38], which was later confirmed in Refs. [40, 45]. The first indication of a narrow resonance within this mass range of the di-photon spectrum was reported at $m_{S}=151.5$ GeV in Ref. [47] and further corroborated with more data in Ref. [48]. [^2] At the time of the manuscript submission new excesses at the same mass have been reported in Ref. [49]. Furthermore, an excess around $95$ GeV, denoted $S^{\prime}$ (which acts as the main source of $b$ quarks), has been observed [50, 51, 52, 53, 44, 48]. This provides a compelling reason to seek Higgs-like resonances near the electroweak (EW) scale, an area where SM backgrounds are quite significant. The use of topological constraints through semi-supervision could be crucial in this context, as they help to dampen the impact of these backgrounds, while reducing model dependencies in the search. The $Z\gamma$ final state dataset is therefore selected as an excellent illustrative example for the proposed methodology.
@@ 2.1 Monte Carlo Simulation
The dominant background for our $S(150\,\mathrm{GeV})\rightarrow Z\gamma$ search is the production of a prompt photon in association with a $Z$ boson [54], comprising over $92\%$ of the total background yield [55]. Our simulation accounts for the fact that the $\ell^{+}\ell^{-}\gamma$ invariant mass of 130 GeV to 170 GeV necessitates the $Z$ boson to be off-shell. We configured MadGraph5 aMC@NLO 2.6.7 to include off-shell $Z$ boson contributions, ensuring a comprehensive representation of the physics involved, although virtual-photon contributions were not considered in our simulation.
The event generation utilized next-to-leading-order (NLO) parton distribution functions combined with NLO matrix elements to ensure consistency and accuracy [56]. Generator-level cuts were applied during the event generation in MadGraph5 aMC@NLO 2.6.7, specifically imposing an invariant mass cut on the $\ell\ell\gamma$ system to select events within the 130-170 GeV range. These generated events were then processed through PYTHIA 8.2 [57] for parton showering and Delphes 3 [58] for detector simulation, where reconstructed-level cuts were applied to mimic experimental conditions. Reconstructed-level cuts included detailed pre-selection criteria such as the number of leptons and photons, their transverse momenta, and isolation requirements. Specifically, the $\ell^{+}\ell^{-}\gamma$ candidates were selected by requiring exactly two oppositely charged same-flavor leptons with an invariant mass greater than 40 GeV and one isolated photon with $E_{T}^{\gamma}>25$ GeV.
Jets were reconstructed using the anti-$k_{t}$ algorithm [59] with the radius parameter $R=0.4$, as implemented in the FastJet 3.2.2 [60] package, and jets with $p_{\textrm{T}}>30$ GeV and $|\eta|<4.7$ are considered. Reconstructed jets overlapping with photons, electrons, or muons in a cone of size $R=0.4$ are removed as well as photons that were within a $\Delta R<0.4$ of any selected lepton. Electrons and muons are required to have $p_{\textrm{T}}>25$ GeV and $|\eta|<2.5$. The central jets are those with a pseudorapidity within the range $|\eta|<2.5$. This cut is applied to ensure that the jets fall within the region of the detector where the tracking and calorimeter systems have optimal performance. Overlap removal procedures were implemented to avoid double-counting objects.
A sample of $000110^{\prime}$ 000 SM $Z\gamma$ MC events are simulated for this analysis. The observables to be used as kinematic features in our study are:
@@
  • The number of jets ($N_{j}$), and the number of central jets ($N_{cj}$.
The feature distributions of the MC dataset are shown in Figure 1. All observables, excluding $m_{\ell\ell\gamma}$, are used as inputs for the NN classifier. The $m_{\ell\ell\gamma}$ distribution, and any features with high correlation with it, cannot be used to train the NN due to the fact that the training samples, mass window ($156144<m_{\ell\ell\gamma}<$ 156 GeV) and side-band ($168132<m_{\ell\ell\gamma}<$ 168 GeV excluding mass window), are defined on this mass.
@@ 3 Methodology
Figure 4: Diagram illustrating the methodology for a pseudo-experiment to assess possible trails factors when using a semi-supervised NN in the search for a narrow resonance. The process starts with the generation of a sample dataset, followed by the training of the semi-supervised NN. Subsequent stages include a background rejection (BR) scan to extract local BR samples from the NN response, fitting signal and background models to each local sample, and finally the computation of the signal significance.
@@ 3.3 Significance calculation for a resonance in the invariant mass distributions
In this analysis, the resonant anomaly detection search strategy is used to discover evidence for BSM physics. The new physics signal is expected to originate from a decay of a new particle, $S$, in the form of a resonance in the corresponding invariant mass. The investigation therefore involves searching for an excess with respect to the continuous SM background. Since we want to examine the probability of a fake signal, we define a signal region (or mass window) in the invariant mass, $m_{\ell\ell\gamma}$, where an excess of signal events can be observed. We select a fixed center of the signal mass window of $m_{\ell\ell\gamma}\approx 150$ GeV, motivated by the corresponding excess [38, 45, 47, 48]. The background (or side-band) region is defined as the mass range surrounding the signal region where only SM background events are expected. The signal (mass-window) range is defined as $144\leq m_{\ell\ell\gamma}\leq 156$ GeV and the background (side-band) range as $168132<m_{\ell\ell\gamma}<$ 168 GeV excluding the mass-window.
@@
where $c_{0}$, $c_{1}$ and $c_{2}$ are fit parameters and $\epsilon_{m}$ is the invariant mass, $m_{\ell\ell\gamma}$, divided by the center of the signal mass window energy, i.e. $13$ TeV. Excesses of signal events within the mass window are captured using a Gaussian function, Equation 2:
$g(m_{\ell\ell\gamma})=c_{s}\cdot e^{-\frac{(m_{\ell\ell\gamma}-\mu_{s})^{2}}{2\cdot\sigma_{s}^{2}}}$ (2)
where $c_{s}$ is the parameter measuring the height of the peak, $\mu_{s}$ is the position of the center of the peak, and $\sigma_{s}$ is the search resolution. In this analysis, the $\mu_{s}$ is fixed at $150$ GeV and $\sigma_{s}$ is set to the search resolution of $2.4$ GeV [55].
@@
where $N_{\rm s}$ and $N_{\rm b}$ are the number of signal and background events respectively. [^4] Note that the systematics related to the background fitting functions, i.e. the spurious signal analysis, are not included here. However, this uncertainty should be included on top of the additional look-elsewhere effect studied here and is not related to the use of NNs. Each pseudo-experiment will therefore produce six local significance values, reflecting each background rejection sample. An example of the signal and background fits for a single iteration of the pseudo-experiment are shown in Appendix A, Figure 8.
@@ 4 Results and Discussion
This methodology, similar to mass scans in high-energy physics, segments our analysis through the use of background rejection (BR) samples, where each threshold acts as a local category. Each category, however, is not statistically independent as the events in each sample overlap. This approach facilitates an examination of the impact of background rejection levels on our results, enabling a nuanced understanding of their influence on our findings. The probability density functions (PDF) of the significances for the six different background rejection cuts ($f(Z)_{\rm BR}$) are shown in Figure 5(a). Note that these distributions have approximately a Gaussian shape, however with a standard deviation between $0.65$ and $0.71$. This decrease in variance w.r.t. the standard Gaussian is due to the constraints within the resonance fit imposed a priori via the shape of the resonance and the side-bands definitions. Note that the peak of the PDFs shifts to the right with an increasing background rejection cut. Such a shift emerges from the interplay between signal and background modelling, amplified by the diminishing event counts in samples subjected to higher background rejection cuts.
@@
where BR is a label indicating the local background rejection cut. We then compute the probability that, for each background reject cut, a significance of up-to a value $Z^{T}=1\sigma$, $2\sigma$ and $3\sigma$ is obtained via the cumulative distribution function (CDF):
@@
The distribution of “global” outcomes, is therefore directly calculated as the ensemble of outcomes across all background rejections. This means that the global distribution includes all significance values from all background rejections. It is shown in Figure 6 from where one can see that it approximates a Gaussian distribution with a standard deviation of $0.69$ centred around a mean of $0.15$. The corresponding global $p$-value is calculated as
@@
Figure 6: Global distribution of the significances, i.e. inclusive case over all the background rejection cuts.
@@
Figure 7: The ratio $p({\rm X})/p({\rm BR}=0\%)$ for $X=10\%,20\%,30\%,40\%,50\%$ BR cuts and $p_{\rm global}/p({\rm BR}=0\%)$, corresponding to the trails factor, as a function of the significance threshold, $Z^{T}$.